# An Accelerated Hybrid Low-Rank Augmented Lagrangian Method for Large-Scale Semidefinite Programming on GPU

### v1

@@ 2 Overview of HALLaR
Moreover, the iterate $\tilde{Y}$ computed in equation (11) has the property that $\tilde{Y}\tilde{Y}^{\top}$ is equal to the usual FW iterate $\alpha yy^{T}+(1-\alpha)YY^{T}$ in the $X$-space; hence $\tilde{Y}$ is an $U$-factor for the Frank-Wolfe iterate in the $X$-space.
3 Accelerating Hallar
In this section, we discuss how to translate the initial CPU implementation of HALLaR efficiently to a scalable and well-optimized GPU implementation.
3.1 GPU architecture, threads, and minimizing CPU/GPU communication
GPUs execute threads in a Single Instruction, Multiple Data (SIMD) fashion: groups of 32 threads, called warps, run the same instruction in lockstep. These warps are organized into blocks, each mapped onto a streaming multiprocessor (SM). Within a block, threads can coordinate via barrier synchronizations (e.g. “__syncthreads”) and share data through fast on‑chip shared memory. Kernels (functions designed to run in parallel in the GPU) launch a grid of blocks to cover the entire data set, so that thousands of warps can run concurrently, hiding memory latency and maximizing arithmetic throughput.
The GPU memory hierarchy further shapes performance. Global memory, while large, has high latency and must be accessed coalesced to achieve high bandwidth. Shared memory resides on each SM and offers very low latency when threads within a block reuse it. Finally, each thread has its own registers for the fastest access; spilling to local memory incurs significant cost. Effective GPU kernels minimize global loads/stores, exploit shared memory for data reuse, and balance register usage to avoid spilling.
Figure 1, adapted from the NVIDIA documentation [48], displays the NVIDIA Hopper architecture used in modern GPUs like the H200. The SMs are the main units of execution inside the GPU. They are grouped into the TPC (Texture Processing Clusters) which are in turn grouped into GPC (GPU Processing Clusters). This hierarchy of clusters allows the GPU to improve the management and distribution of the workload between the SMs, improving scalability. The number of available GPCs and TPCs depends on the GPU model. Inside each SM, the warp scheduler, dispatch unit and the L0 instruction cache work as a control unit, delivering instructions to the warps. Next, we see the set of registers each SM has access to, followed by execution components: CUDA cores, tensor cores, LD/ST and SFU. The CUDA cores execute scalar aritmetic operations, while the tensor cores are specialized to execute multiplication and addition $C=D*A+C$ in batches of data. The LD/ST units performs load and store operations on the data. The SFU (Special Function Units) are specialized for transcendental and trigonometric operations. The TMA (Tensor Memory Acceleration) provides asynchronous transfers of large blocks of data between the global and shared memory. Next, we see the L1-cache that optimizes transfers between the shared memory and the CUDA/tensor cores. Back to the GPU, we have the L2-cache that optimizes transfers from the global GPU memory to all SMs. Access to L2-cache by the SMs is controlled by the GPC.
Figure 1: GPU architecture, adapted from the NVIDIA documentation [48], displaying the general GPU design and the components of each SM.
Before launching a kernel, we must transfer the input dataset from the CPU (also known as host in GPU programming) to the GPU (also called device) and define the number of blocks and threads used. The memory allocated in the CPU is not accessible by the GPU and vice-versa. When the kernel is launched, the workload is distributed among the GPCs that, in turn, distribute the workload among their TPCs. The TPCs forward the thread blocks among its SMs. The SMs divide the thread blocks into warps and prepare the shared memory and registers. The warp scheduler and the dispatch unit select the warps ready to run and dispatch the instructions to execution by the CUDA/tensor cores, the LD/ST or the SFU. After the kernel is finished, a transfer from the device back to the host may be necessary.
The GPU programming paradigm is meant to integrate CPU and GPU to obtain the performance gains of both architectures. Sections of the code that allow high parallelization, like matrix, vector and tensor operations are implemented as kernels, while serial tasks are implemented as regular functions in the CPU. This CPU/GPU integration sometimes requires data transfers between the CPU and GPU. Modern GPUs compute faster than they can feed data from memory, so these transfers are often the bottleneck of applications. Therefore we must minimize data transfers to achieve maximum performance.
cuHALLaR stores on the GPU, at the beginning, key data such as the primal/dual solutions, the right-hand side vector $b$, and the data used to compute $C$ and $\mathcal{A}$. Any other auxiliary data is allocated directly on the GPU. So all operations that involve arrays with huge size are done in the GPU, including the dual feasibility computations, which was computed in the CPU by other GPU solvers like [30].
@@ 3.1 GPU architecture, threads, and minimizing CPU/GPU communication
Another important performance aspect is that, just like in the CPU, launching threads in the GPU incurs an overhead due to the grid, context and communication management. Therefore it is also important to minimize the number of kernels launched, merging trivial kernels and avoiding to launch them inside loops when possible.
In HALLaR, three operations are critical to performance: the linear map $\mathcal{A}$, the adjoint $\mathcal{A}^{*}$ and the gradient $\nabla g$. These operations are performed at every ADAP-FISTA iteration, and the gradient is also computed at the Frank-Wolfe step of HLR. For the maximum stable set and matrix completion problems, CuHallar solves $\mathcal{A}$ and $\mathcal{A}^{*}$ in a single kernel, and the gradient is solved with two kernels. In the phase retrieval problem, due to the computation of the discrete Fourier transform, all operations are solved with $r$ kernel calls, where $r$ is the rank of the current solution.
In cuHALLaR, we use the CUDA.jl library, the Julia interface for the CUDA runtime library. CUDA.jl provides high level implementations of many cuBLAS, cuSPARSE and cuFFT functions. The transfer of data between the CPU and GPU is as simple as converting a variable between the Julia Array type and the CUDA.jl CuArray type. CUDA.jl also provides macros so users can implement their own kernels. With the help of Julia broadcast fusion feature, CUDA.jl is able to merge kernels in broadcasted expressions. This allows cuHallar to reduce the number of kernels invoked, specially in the ADAP-FISTA method.

### v2

@@ 2 Overview of HALLaR
Moreover, the iterate $\tilde{Y}$ computed in equation (12) has the property that $\tilde{Y}\tilde{Y}^{\top}$ is equal to the usual Frank-Wolfe iterate $(1-\alpha)YY^{\top}+\alpha(\tau^{2}yy^{\top})$ in the $X$-space; hence $\tilde{Y}$ is a valid $U$-factor for the Frank-Wolfe iterate.
3 Computational Aspects of cuHALLaR
GPU architecture is designed to perform many operations through massive parallelism, using a single instruction multiple data (SIMD) model. Unlike CPUs, which have fewer cores and are optimized for sequential execution and control, GPUs have thousands of simpler cores that are organized in streaming multiprocessors (SMs). GPUs possess a memory hierarchy similar to CPUs with fast and small registers for each SM, a shared memory between blocks of SMs, and a slower and bigger global GPU memory.
In this section, we discuss the key computational factors that allow cuHALLaR to achieve significant speedups over HALLaR through the use of GPU programming. This section is divided into two subsections. The first subsection discusses effectively dividing tasks between the CPU and the GPU. The second subsection describes our new way of efficiently reading and inputting SDP data that is suitable for GPU programming.
3.1 Effective Division of CPU and GPU Tasks
The primary computational concern in GPU programming is effectively dividing tasks between the CPU and GPU to exploit the advantages of both architectures. Highly parallelizable operations, such as matrix, vector and tensor operations should be implemented in the GPU as functions which are called kernels. On the other hand, serial tasks should be implemented as regular functions in the CPU. Hence, the primary computational bottleneck in modern GPUs is the possibly large number of data transfers that need to be performed between the CPU and the GPU.
@@ 3.1 Effective Division of CPU and GPU Tasks
Another important computational consideration in GPU programming is that, like in the CPU, launching threads in the GPU also incurs an overhead due to thread management and communication. Therefore, an efficient GPU implementation must minimize the number of data transfers and the number of kernels launched.
To minimize CPU and GPU communication overhead, cuHALLaR pre-allocates essential problem data from (P) on the GPU, including the objective matrix $C$, the linear operator $\mathcal{A}$, and the right-hand side vector $b$. All subsequent computations involving large matrices and vectors are performed exclusively on the GPU.
3.2 Reading Data Efficiently and Input Format
SDPA format.
Introduced in [74, 75, 26], the SDPA format allows for the specification of SDPs using sparse matrix representations of the cost and constraint matrices. Each matrix is stored by explicitly listing all nonzero entries in its upper triangular part, which is efficient for problems where all data matrices are sparse. However, for SDP instances with a dense but well-structured data matrix (e.g., the matrix of all ones or the identity matrix plus a low-rank matrix), it becomes prohibitively expensive to store and handle their input data in SDPA format. For example, if $C$ is the matrix of all ones, then SDPA format requires the storage of $n(n+1)/2$ entries, and hence does not take advantage of the fact that such a matrix can be represented by either a single scalar or a single $n$-vector.
HSLR format.
To address the issues of the SDPA format and to improve handling of large-scale problems, we introduce a new input format, namely, Hybrid Sparse Low-Rank (HSLR) format, where each constraint matrix $A_{\ell}$ (with the convention $C=A_{0}$) can be decomposed as:
$A_{\ell}=A_{\ell}^{\mathrm{sp}}+A_{\ell}^{\mathrm{lr}},\qquad A_{\ell}^{\mathrm{lr}}=K_{\ell}D_{\ell}K_{\ell}^{\top},\qquad\ell=0,1,\dots,m$ (18)
where $A_{\ell}^{\mathrm{sp}}\in\mathbb{S}^{n}$ is a sparse symmetric matrix, $K_{\ell}\in\mathbb{R}^{n\times r_{\ell}}$, and $D_{\ell}\in\mathbb{S}^{r_{\ell}}$ is a symmetric matrix that is not necessarily diagonal. This structure allows us to reformulate the core operations of cuHALLaR in terms of sparse matrix algebra and small dense matrix multiplications. In particular, cuHALLaR never has to form any $n\times n$ matrices. For examples on how to construct HSLR format for several structured SDP instances such as the SDP relaxations of the matrix completion and maximum stable set problems, the reader should refer to our user manual [1].
The advantages of HSLR over sparse SDPA format are twofold:
  1. 1.
    Storage Efficiency: HSLR format has a significantly lower memory footprint than SDPA format. Specifically, if a data matrix $A_{\ell}$ is such that $A_{\ell}^{\mathrm{l}r}\neq 0$ and $D_{\ell}$ is diagonal, then it requires $n_{\ell}+r_{\ell}n+r_{\ell}$ entries (resp., n(n+1)/2) to be stored in HSLR (resp., SDPA) format, where $n_{\ell}$ is the number of nonzeros of $A_{\ell}^{\mathrm{s}p}$.
  2. 2.
    Computational Efficiency: During cuHALLaR’s execution, matrix-vector products and trace evaluations are computed by separately handling the sparse and low-rank components of each data matrix $A_{\ell}$. Since cuHALLaR handles these components separately, it never has to form the possibly dense matrix $A_{\ell}$ and calls different kernels based on the specific structure of each $A_{\ell}$.
4 Efficient Implementation of $\tilde{\mathcal{A}}(UU^{\top})$ and $\tilde{\mathcal{A}}^{*}(\tilde{p})U$
The efficient evaluations of the linear maps $\mathcal{A}(X):\mathbb{S}^{n}\rightarrow\mathbb{R}^{m}$ and $\mathcal{A}^{*}(p):\mathbb{R}^{m}\rightarrow\mathbb{S}^{n}$ as in (1) are essential to the performance of cuHALLaR. To avoid the formation of a possibly dense $n\times n$ matrix $X$, cuHALLaR forms a low-rank factor $U\in\mathbb{R}^{n\times r}$ of $X$ that satisfies $X=UU^{\top}$. At every iteration, cuHALLaR has to evaluate for a given $(U,p)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m}$ the quantities $\tilde{{\cal A}}(UU^{\top})$ and $\tilde{{\cal A}}^{*}(\tilde{p})U$ where $\tilde{p}=(1,p)\in\mathbb{R}^{m+1}$, and $\tilde{\mathcal{A}}:\mathbb{S}^{n}\rightarrow\mathbb{R}^{m+1}$ and $\tilde{{\cal A}}^{*}:\mathbb{R}^{m+1}\to\mathbb{S}^{n}$ are given by
$\tilde{\mathcal{A}}(X):=\begin{bmatrix}A_{0}\bullet X\\ A_{1}\bullet X\\ \vdots\\ A_{m}\bullet X\end{bmatrix}:=\begin{bmatrix}A_{0}\bullet X\\ \mathcal{A}(X)\end{bmatrix}\,,\qquad\tilde{\mathcal{A}}^{*}(\tilde{p}):=\sum_{\ell=0}^{m}\tilde{p}_{\ell}A_{\ell}=A_{0}+\sum_{\ell=1}^{m}p_{\ell}A_{\ell}=A_{0}+{\mathcal{A}}^{*}(p)$ (19)
where $A_{0}:=C$ and $\mathcal{A}(\cdot)$ and $\mathcal{A}^{*}(\cdot)$ are as in (1). It follows from the definitions of $\nabla_{U}\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})$ and $G$ in (15) and (16), respectively, that cuHALLaR has to evaluate $\tilde{\mathcal{A}}(UU^{\top})$ in steps 6, 7, 11, and 14, and $\tilde{\mathcal{A}}^{*}(\tilde{p})U$ in steps 6 and 11.
This section presents the details of the efficient implementation of the operations in (19) in the GPU, using the fact that the data matrices $A_{\ell}$ are assumed to be the sum of a sparse and a low-rank matrix as in (18).
4.1 Constraint Evaluation $\tilde{\mathcal{A}}(UU^{\top})$
[…]
$T^{\mathrm{s}p}_{\ell}=A^{\mathrm{s}p}_{\ell}\bullet(UU^{\top}),\quad T_{\ell}^{\mathrm{l}r}=A^{\mathrm{l}r}_{\ell}\bullet(UU^{\top}),\quad\ell=0,1,\dots,m.$ (20)
In practice, cuHALLaR implements efficient subroutines to separately compute $T^{\mathrm{s}p}$ and $T^{\mathrm{l}r}$. The rest of this subsection discusses in more detail how cuHALLaR computes these two quantities efficiently.
Sparse Component.
This paragraph describes how cuHALLaR efficiently computes $T^{\mathrm{s}p}\in\mathbb{R}^{m+1}$. Let $B=UU^{\top}$. It is easy to see that the entries of $B$ can be computed as $B_{ij}=\langle U_{i,:},U_{j,:}\rangle$, where $U_{i,:}$ and $U_{j,:}$ are the $i$-th and $j$-th rows of $U$, respectively. The $\ell$-th component of $T^{\mathrm{sp}}$, $T_{\ell}^{\mathrm{sp}}=A_{\ell}^{\mathrm{sp}}\bullet(UU^{\top})$, is then computed as:
$T_{\ell}^{\mathrm{sp}}=\sum_{i=1}^{n}\sum_{j=1}^{n}(A_{\ell}^{\mathrm{sp}})_{ij}B_{ij}.$ (21)
Since $A_{\ell}^{\mathrm{sp}}$ and $B$ are symmetric matrices, this summation can be rewritten by defining the weights
$\omega_{ij,\ell}\vcentcolon=(2-\delta_{ij})(A_{\ell}^{\mathrm{sp}})_{ij},\text{ where }\delta_{ij}=\begin{cases}1\quad\text{if }i=j\\ 0\quad\text{otherwise}.\end{cases}$ (22)
Using (22), we see that $T_{\ell}^{\mathrm{sp}}=\sum_{i\leq j}\omega_{ij,\ell}B_{ij}$.
To avoid redundant computations of $B_{ij}$ across different constraints $\ell$, we reorganize this computation by using the union of the supports of $A_{\ell}^{\mathrm{s}p}$. This union $\mathcal{S}$ is formally defined as:
$\mathcal{S}\vcentcolon=\bigcup_{\ell=0}^{m}\mathrm{supp}(A_{\ell}^{\mathrm{sp}})\cap\{(i,j):1\leq i\leq j\leq n\}.$ (23)
For each pair $(i,j)\in\mathcal{S}$, we define $\mathcal{L}_{ij}$ to be the set of indices $\ell$ for which the entry $(i,j)$ is active:
$\mathcal{L}_{ij}\vcentcolon=\{\ell:(i,j)\in\mathrm{supp}(A_{\ell}^{\mathrm{sp}})\}.$ (24)
To compute $T^{\mathrm{s}p}$, we then compute the quantity $B_{ij}$ over each pair $(i,j)\in\mathcal{S}$ and update the corresponding entries $T^{\mathrm{s}p}_{\ell}$ for all $\ell\in\mathcal{L}_{ij}$. This procedure is formally described in Algorithm 2 below.
[…]
1:Input: $U\in\mathbb{R}^{n\times r}$, data matrices $\{A_{\ell}^{\mathrm{sp}}\}_{\ell=0}^{m}\subseteq\mathbb{S}^{n}$, and precomputed structures $\mathcal{S}$ and $\{\mathcal{L}_{ij}\}$ from (23)–(24).
2:Output: $T^{\mathrm{s}p}\in\mathbb{R}^{m+1}$ where $T^{\mathrm{s}p}_{\ell}=A^{\mathrm{s}p}_{\ell}\bullet UU^{\top}$.
3:Initialize $T^{\mathrm{s}p}\leftarrow 0\in\mathbb{R}^{m+1}$.
4:for all $(i,j)\in\mathcal{S}$ in parallel do
5:  $B_{ij}\leftarrow\langle U_{i,:},U_{j,:}\rangle$
6:  for all $\ell\in\mathcal{L}_{ij}$ do
7:   $\omega_{ij,\ell}\leftarrow(2-\delta_{ij})(A_{\ell}^{\mathrm{sp}})_{ij}$
8:   $T^{\mathrm{s}p}_{\ell}\leftarrow T^{\mathrm{s}p}_{\ell}+\omega_{ij,\ell}B_{ij}$
9:  end for
10:end for
11:return $T^{\mathrm{s}p}$
In Algorithm 2, the outer loop over the set $\mathcal{S}$ is parallelized. More specifically, each unique index pair $(i,j)\in\mathcal{S}$ is assigned to a separate GPU thread. Each thread performs a loop over $r\ll n$ to compute $B_{ij}$ and a loop over $|\mathcal{L}_{ij}|\ll m+1$ that updates $T^{\mathrm{s}p}_{\ell}$.
Low-Rank Component.
This paragraph describes how cuHALLaR efficiently computes $T^{\mathrm{l}r}\in\mathbb{R}^{m+1}$. Using the second relation in (18), the $\ell$-th component of $T^{\mathrm{lr}}$, $T_{\ell}^{\mathrm{lr}}=A_{\ell}^{\mathrm{lr}}\bullet(UU^{\top})$, can be computed as
$T_{\ell}^{\mathrm{lr}}=\mathrm{Tr}(K_{\ell}D_{\ell}K_{\ell}^{\top}(UU^{\top}))=\mathrm{Tr}(D_{\ell}(K_{\ell}^{\top}U)(U^{\top}K_{\ell})).$ (25)
Defining $Y_{\ell}\vcentcolon=K_{\ell}^{\top}U\in\mathbb{R}^{r_{\ell}\times r}$, relation (25) further simplifies to:
$T_{\ell}^{\mathrm{lr}}=\mathrm{Tr}(D_{\ell}Y_{\ell}Y_{\ell}^{\top}).$ (26)
Algorithm 3 below formally describes our efficient procedure for computing $T^{\mathrm{lr}}$.
[…]
1:Input: $U\in\mathbb{R}^{n\times r}$ and problem data $\{K_{\ell},D_{\ell}\}_{\ell=0}^{m}$.
2:Output: $T^{\mathrm{l}r}\in\mathbb{R}^{m+1}$ where $T^{\mathrm{l}r}_{\ell}=A^{\mathrm{l}r}_{\ell}\bullet UU^{\top}$.
3:Initialize $T^{\mathrm{l}r}\leftarrow 0\in\mathbb{R}^{m+1}$.
4:for $\ell=0$ to $m$ do
5:  $Y_{\ell}\leftarrow K_{\ell}^{\top}U$ $\triangleright$ in parallel
6:  $T^{\mathrm{l}r}_{\ell}\leftarrow 0+\mathrm{Tr}(D_{\ell}Y_{\ell}Y_{\ell}^{\top})$
7:end for
8:return $T^{\mathrm{l}r}$
In Algorithm 3, the loop over $\ell=0,\dots,m$ is performed in serial, while the computations of $Y_{\ell}$ and $\mathrm{Tr}(D_{\ell}Y_{\ell}Y_{\ell}^{\top})$ are performed in parallel. In our numerical experiments, we found that parallelizing these computations was very efficient.
4.2 Adjoint $\tilde{\mathcal{A}}^{*}(\tilde{p})U$
This subsection details how, for a given $(U,p)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m}$, cuHALLaR efficiently computes the quantity $W=\tilde{\mathcal{A}}^{*}(\tilde{p})U$ as in (19), where $\tilde{p}=(1,p)\in\mathbb{R}^{m+1}$. It follows from (18) and the second identity in (19) that $W$ is given by
$W=\underbrace{\left(\sum_{\ell=0}^{m}\tilde{p}_{\ell}A_{\ell}^{\mathrm{sp}}\right)U}_{W^{\mathrm{sp}}}+\underbrace{\sum_{\ell=0}^{m}\tilde{p}_{\ell}(K_{\ell}D_{\ell}K_{\ell}^{\top})U}_{W^{\mathrm{lr}}}\,.$ (27)
The remaining part of this subsection discusses how cuHALLaR implements efficient subroutines to compute $W^{\mathrm{s}p}$ and $W^{\mathrm{l}r}$ separately.
Sparse Component.
To compute $W^{\mathrm{s}p}$ as in (27), we first allocate a sparse symmetric matrix $M$ with nonzero entries at indices $(i,j)$, where $\mathcal{L}_{ij}$ in (24) is nonempty. After doing this allocation, we then update the values of the nonzero entries of $M_{ij}$ in parallel using the formula:
$M_{ij}=\sum\limits_{\ell\in\mathcal{L}_{ij}}\tilde{p}_{\ell}(A^{\mathrm{sp}}_{\ell})_{ij}.$ (28)
The matrix $W^{\mathrm{sp}}$ is then computed as $W^{\mathrm{sp}}=MU$. This procedure is formally described in Algorithm 4 below.
Algorithm 4 Computation of $W^{\mathrm{s}p}$ as in (27).
1:Input: $U\in\mathbb{R}^{n\times r}$, $\tilde{p}=(1,p)\in\mathbb{R}^{m+1}$ where $p\in\mathbb{R}^{m}$, $\{A_{\ell}^{\mathrm{sp}}\}_{\ell=0}^{m}\subseteq\mathbb{S}^{n}$, and precomputed structures $\mathcal{S}$, $\{\mathcal{L}_{ij}\}$ from (23)–(24).
2:Output: $W^{\mathrm{s}p}\in\mathbb{R}^{n\times r}$ where $W^{\mathrm{s}p}=\left(\sum_{\ell=0}^{m}\tilde{p}_{\ell}A_{\ell}^{\mathrm{sp}}\right)U$.
3:Initialize $M\in\mathbb{S}^{n}$ with pattern $\mathcal{S}\cup\{(j,i):(i,j)\in\mathcal{S}\}$.
4:for all $(i,j)\in\mathcal{S}$, in parallel do
5:  $w\leftarrow 0$
6:  for all $\ell\in\mathcal{L}_{ij}$ do
7:   $w\leftarrow w+\tilde{p}_{\ell}(A^{\mathrm{sp}}_{\ell})_{ij}$
8:  end for
9:  $M_{ij}\leftarrow w$
10:  if $i\neq j$ then
11:   $M_{ji}\leftarrow w$
12:  end if
13:end for
[…]
15:return $W^{\mathrm{s}p}$
Several remarks about Algorithm 4 are now given. The matrix $M$, which is initialized in step 3, is stored in sparse format with symmetric pattern $\mathcal{S}\cup\{(j,i):(i,j)\in\mathcal{S}\}$, hence it contains at most $2|\mathcal{S}|-|\{\,i:(i,i)\in\mathcal{S}\,\}|$ nonzero entries. Also, it is worthwhile to note that the outer for loop over $\mathcal{S}$ is parallelized. More specifically, each unique index pair $(i,j)\in\mathcal{S}$ is assigned to a separate GPU thread. Each thread then performs a loop over $|\mathcal{L}_{ij}|\ll m+1$ to compute $w=\sum_{\ell\in\mathcal{L}_{ij}}\tilde{p}_{\ell}(A^{\mathrm{sp}}_{\ell})_{ij}$ and updates the corresponding entries in $M$.
Low-Rank Component.
This paragraph describes how cuHALLaR efficiently computes $W^{\mathrm{l}r}\in\mathbb{R}^{n\times r}$. Using the second quantity in (27) and defining $Y_{\ell}=K_{\ell}^{\top}U\in\mathbb{R}^{r_{\ell}\times r}$, it follows that $W^{\mathrm{lr}}$ can be computed as
$W^{\mathrm{lr}}=\sum_{\ell=0}^{m}\tilde{p}_{\ell}K_{\ell}(D_{\ell}Y_{\ell}).$ (29)
Algorithm 5 below formally describes our efficient procedure for computing $W^{\mathrm{lr}}$.
Algorithm 5 Computation of $W^{\mathrm{l}r}$ as in (27).
1:Input: $U\in\mathbb{R}^{n\times r}$, $\tilde{p}=(1,p)\in\mathbb{R}^{m+1}$ where $p\in\mathbb{R}^{m}$, and problem data $\{K_{\ell},D_{\ell}\}_{\ell=0}^{m}$.
2:Output: $W^{\mathrm{l}r}\in\mathbb{R}^{n\times r}$ where $W^{\mathrm{l}r}=\left(\sum_{\ell=0}^{m}\tilde{p}_{\ell}A_{\ell}^{\mathrm{lr}}\right)U$.
3:Initialize $W^{\mathrm{l}r}\leftarrow 0_{n\times r}$.
4:for $\ell=0$ to $m$ do
5:  $Y_{\ell}\leftarrow K_{\ell}^{\top}U$
6:  $Z_{\ell}\leftarrow D_{\ell}Y_{\ell}$ $\triangleright$ in parallel
7:  $W^{\textrm{l}r}\leftarrow W^{\textrm{l}r}+\tilde{p}_{\ell}(K_{\ell}Z_{\ell})$
8:end for
9:return $W^{\textrm{l}r}$
In Algorithm 5, the loop over $\ell=0,\dots,m$ is performed in serial, while the computations of $Y_{\ell}=K_{\ell}^{\top}U$ and $Z_{\ell}=D_{\ell}Y_{\ell}$ used to update $W^{\mathrm{l}r}$ are performed in parallel.
4.3 Kernels for problems in SDPA format
This subsection describes cuHALLaR’s efficient approach for dealing with SDP instances given in sparse SDPA format. The parser reads each data matrix $A_{\ell}$, for $\ell=0,...,m$ and computes its density as
$d_{\ell}=\frac{nnz(A_{\ell})}{n^{2}}.$ (30)
The kernels for SDPA format separate constraints into sparse or dense. Let $A_{\ell}^{dn}$ be the representation of data matrix $A_{\ell}$ in dense format. If $d_{\ell}>10^{-1}$, cuHALLaR sets $A_{\ell}=A_{\ell}^{dn}$; otherwise it sets $A_{\ell}=A_{\ell}^{\mathrm{s}p}$. The quantities $\mathcal{S}$ and $\{\mathcal{L}_{ij}\}$ in (23) and (24) are precomputed from the data matrices which are determined to be sparse. The evaluations of $\tilde{\mathcal{A}}(UU^{\top})$ and $\tilde{\mathcal{A}}^{*}(\tilde{p})U$ use similar approaches as the ones described in Section 4. The sparse components of these evaluations are performed using Algorithms 2 and 4 while the dense components of these evaluations are evaluated using adaptations of Algorithms 3 and 5, where $A_{\ell}^{dn}$ replaces $A_{\ell}^{\mathrm{l}r}=K_{\ell}D_{\ell}K_{\ell}^{\top}$.